X^2d^2y/dx^2-3xdy/dx+4y=x+x^2logx
1
Expert's answer
2021-09-20T11:35:33-0400

Let us solve the differential equation "x^2\frac{d^2y}{dx^2}-3x\frac{dy}{dx}+4y=x+x^2\ln x." For this let us use the transformation "x=e^t." Then "y'_x=y'_te^{-t},\ y''_{x^2}=(y''_{t^2}-y'_t)e^{-2t}."


Then we get the following equation

"e^{2t}(y''_{t^2}-y'_t)e^{-2t}-3e^ty'_te^{-t}+4y=e^t+te^{2t}," which is equivalent to "y''_{t^2}-4y'_t+4y=e^t+te^{2t}."


The characteristic equation "k^2-4k+4=0" of the equation "y''_{t^2}-4y'_t+4y=0" is equivalent to "(k-2)^2=0," and hence has the solutions "k_1=k_2=2."


It follows that the general solution of the equation "y''_{t^2}-4y'_t+4y=e^t+te^{2t}" is "y(t)=(C_1+C_2t)e^{2t}+y_p," where "y_p=ae^t+t^2(bt+c)e^{2t}."


Then

"y'_p(t)=ae^t+(3bt^2+2ct)e^{2t}+(bt^3+ct^2)2e^{2t}=\nae^t+(2bt^3+(3b+2c)t^2+2ct)e^{2t},"

"y''_p(t)=ae^t+(6bt^2+2(3b+2c)t+2c)e^{2t}+2(2bt^3+(3b+2c)t^2+2ct)e^{2t}=\\
ae^t+(4bt^3+(12b+4c)t^2+(6b+6c)t+2c)e^{2t}."


Then we get "ae^t+(4bt^3+(12b+4c)t^2+(6b+6c)t+2c)e^{2t}-4(ae^t+(2bt^3+(3b+2c)t^2+2ct)e^{2t})+4(ae^t+(bt^3+ct^2)e^{2t})=e^t+te^{2t}."


It follows that

"ae^t+((6b-2c)t+2c)e^{2t}=e^t+te^{2t}."


Then we get "a=1,\ 6b-2c=1,\ 2c=0," and hence "a=1,\ b=\frac{1}6,\ c=0."


Therefore, the solution of "y''_{t^2}-4y'_t+4y=e^t+te^{2t}" is

"y(t)=(C_1+C_2t)e^{2t}+e^t+\frac{1}6t^3e^{2t}."


Consequently, the solution of "x^2\frac{d^2y}{dx^2}-3x\frac{dy}{dx}+4y=x+x^2\ln x" is


"y(x)=(C_1+C_2\ln x)x^2+x+\frac{1}6x^2\ln^3x."